A health care provider has ordered dobutamine 10mcg/kg/min for your 68 YO patient who weighs 172...

  1. A health care provider has ordered dobutamine 10mcg/kg/min for your 68 YO patient who weighs 172 lbs. The dobutamine is available as 1000mg/250mL. At what rate will you set the IV infusion pump?
Answer #1

In order to calculate this first weight has to be converted to kg ,then disease of drug to be calculated, then minutes to be calculated, finally substituted in the infusion formula (volume(ml÷ minutes×drop factor)

Given

Order:Dobutamine

Weight:172lbs ( 0.454×172=78kg)

Dose:10mcg/kg/min

When 1 kg =10mcg

Then 78 kg=??McGee

Cross multiply and divide

=78×10÷1

=78×10

=780mcg

So in a minute 780mcg will be administered

Available:

250ml/1000mg (1000mg=1000×1000=1000000mcg)

When

1minute=780mcg

So 60 minute (1 hr)=???

Cross multiply and divide

=60×780÷1

=60×780

=46800mcg

So in 1 hr 46800 mcg drug will be administered

Therefore 1000000 mg will be administered in what time

When 46800mcg=60min

Then 1000000mcg=??min

Cross multiply and divide

=1000000×60÷46800

=60000000÷46800

=1282 min

Flow rate: Volume(ml)÷ minutes×drop factor

Here 15 gtt/ml is the standard drop factor

=250÷1282×15

=0.19×15

=2.9 or 3 gtt/min

The infusion pump has to be set at 3 gtt/min